# Recent advances in text embedding: A Comprehensive Review of Top-Performing Methods on the MTEB Benchmark

### v1

@@ 5 LLMs focused universal text embeddings
LLMs are are extensively pre-trained on diverse large quantity of web-scale data, which can be used to improve the universal text embeddings in two different ways as summarized in Table 2. Firstly, LLMs can be used to generate high quality multilingual multi-task synthetic data as demonstrated by researchers from Microsoft and Google (?, ?, ?). Secondly, LLMs can be used as backbone for text embeddings as they do not need the contrastive pre-training step used in existing state of the art text embedding models. For example, E5-mistral-7b-instruct perform multi-task fine-tuning on Mistral 7b model which is one of the best performing method on MTEB. Echo-mistral (?), LLM2Vec (?), gte-Qwen1.5-7B-instruct (?) and GRIT (?) propose various different solutions so that decoder only LLMs with casual attention can generate high quality text embeddings using bidirectional attention. More details about how different universal text embeddings leverage LLMs to improve their universality can be found below.
@@
E5-mistral-7b-instruct is one of the best performing text embeddings on the MTEB benchmark, which is also a representative text embedding model leveraging LLMs. Firstly, proprietary LLMs including GPT-35-Turbo and GPT-4 are used to generate synthetic data covering a diverse range of text embedding tasks in 93 languages (among which 25% are generated by GPT-35-Turbo and others are generated by GPT-4) (?). In terms of the quality generated data, the authors find that the overall quality is acceptable despite a portion of GPT-35-Turbo outputs do not follow the instructions in the prompt templates strictly. Secondly, pre-trained open source LLM Mistral-7b checkpoint (?) is selected to be fine-tuned on a mixture of synthetic and labeled data (collection of 13 public datasets) with around 1.8M examples after sampling. One advantage of using LLMs such as Mistral (?) for text embedding is that they are extensively pre-trained on web-scale data already, which does not need the contrastive pre-training step used in existing state of the art text embedding models. Given a pre-trained LLM, an [EOS] token is appended to the end of the query and document. The last layer [EOS] vector is used as the text embeddings. To help the model accommodate different tasks, instruction templates (which are used by all LLMs focused universal text embeddings described in this section as well as some of the previously mentioned universal text embeddings such as BGE (?) ) are applied to the original query $q^{+}$ to generate a new one $q_{inst}^{+}$ given a relevant query-document pair $(q^{+},d^{+})$:
@@
where “task_definition” is a placeholder for a one-sentence description of the embedding task added only to the query side but not to the document side (?). Some other details about E5-mistral-7b-instruct include:
  • Fine-tuning data: generated synthetic data, ELI5 (?) (sample ratio 0.1), HotpotQA (?), FEVER (?), MIRACL (?), MS-MARCO passage ranking (sample ratio 0.5) and document ranking (sample ratio 0.2) (?), NQ (?), NLI (?), SQuAD (?), TriviaQA (?), Quora Duplicate Questions (?) (sample ratio 0.1), Mr-TyDi (?), DuReader (?), and T2Ranking (?) (sample ratio 0.5) datasets.
@@
  • Negative sampling: in-batch negative samples and hard negatives (for the datasets without hard negatives, mE5base (?) is used to to mine top 100 hard negatives).
  • Model size: 7B (42M trainable parameters using Low-rank adaptation (LoRA) (?))
The experimental results from (?) shows that even with only synthetic data, the performance of E5-mistral-7b-instruct on MTEB English benchmark is still very competitive. E5-mistral-7b-instruct also has the multilingual capabilities with good performances over high-resource languages. Furthermore, the authors discovered that the method of incorporating instructions has a considerable impact on the performance. Their hypothesis is that the model is better informed about the embedding task at hand through natural language instructions, thereby allowing the model to produce more distinctive embeddings (?).
@@
Built on top of the E5-mistral-7b-instruct, SFR-Embedding-Mistral is also one of the top-ranking universal text embeddings on the MTEB English benchmark with 0.93% performance increase compared to E5-mistral-7b-instruct. The authors summarized their main takeaways in (?) (the detailed report is not released) as:
@@
Even though constructing text embeddings from autoregressive pretrained LLMs seems promising, the authors from (?) identified a striking failure mode of autoregressive language models trained on the next-token objective: the contextualized token embeddings, represented by the vector of last-hidden-layer activations at a specific input token’s position, lack information from tokens appearing later in the sentence because of the causal attention mask. Given the following example provided in (?):
@@
In this example, the classical LLMs based contextualized embeddings of the first half of $d^{-}$ and $d^{+}$ are both similar to q because they do not attend to the second half of the sentence, which leads to the overestimation of the similarity between q and $d^{-}$ by any pooling strategy that uses information from the first half (?).
To mitigate this striking failure mode and take advantage of the bidirectional context information, a simple fix is proposed by presenting the input sentence twice to LLMs. The final contextualized embeddings can then be extracted from the second occurrence of the sentence. LLMs are instructed to undertake basic task such as rewriting or repeating in order to prompt the second occurrence to effectively ”encode” information from the first (?). Despite twice the computational cost of classical embeddings, experimental results show that Echo embeddings can improve the LLM based text embedding quality significantly under both zero-shot setting and fine-tuning setting. Some other details about Echo-mistral (echo-mistral-7b-instruct-last) include:
  • Fine-tuning data: same as E5-mistral-7b-instruct (?) without synthetic data
@@
Similar to the idea of Echo-mistral, the authors of (?) believe that the slow adoption of decoder-only LLMs in text embedding tasks is partly due to their causal attention mechanism, which restricts their ability to create bidirectional contextualized representations from encompassing information from the whole input sequence (a necessary trade-off for generative capabilities). Improving the architectural flaw of decoder-only LLMs for text embedding tasks is highly desirable because: 1. decoder-only LLMs are much more sample-efficient than encoder-only models (?); 2. LLMs are supported by a robust ecosystem, including comprehensive tools and well tested pre-training techniques, leading to their continuous enhancement by the community; 3. the good instruction following ability of LLMs (?, ?) makes them ideal for creating universal text embedding models that can handle a wide range of tasks using instructions.
To improve the text embeddings from decoder-only LLMs, LLM2Vec proposes a simple unsupervised approach that can transform any decoder-only LLM into a strong text encoder in three simple steps: 1. enabling bidirectional attention by replacing the causal attention mask of decoder-only LLMs with an all-ones matrix; 2. Masked Next Token Prediction (MNTP): combining next token prediction with masked language modeling (?) to make the model aware of its bidirectional attention; and 3. unsupervised contrastive learning for better sequence representations: the model processes an input sequence twice with independently sampled dropout masks to generate two distinct representations, and is trained to increase the similarity between these representations while decreasing similarity with other sequence representations in the batch (?). Their empirical results show that LLMs can be efficiently converted into universal text embeddings without requiring costly adaptation or synthetic GPT-4 generated data. Some other details about LLM2Vec include:
@@
  • Supervised contrastive learning data: adaptations of E5 (?): the public portion of the E5 dataset (?) curated by (?)
@@
Similar to the idea from Echo-mistral and LLM2Vec, the authors in (?) also highlight the importance of bidirectional attention for general purpose universal text embeddings. However, GRIT takes the general purpose model to the next level by training a large language model to handle both generative and embedding tasks (all text-based language problems) distinguished through instructions.
Both representational instruction tuning (?, ?, ?) and generative instruction tuning (?, ?, ?) are combined into one unified model by GRIT. Firstly, GRIT uses bidirectional attention with mean pooling over the final hidden state to get the text embedding. Contrastive objective with in-batch negatives are used to finetune a pretrained large language model following prior works (?, ?). The average of the final hidden states of only the input sample is calculated during mean pooling, while disregarding the instruction and format tokens. Nonetheless, these tokens continue to impact the final representation via the self-attention mechanism (?). Secondly, the language modeling objective of next token prediction (?) is used to compute the loss on generative data, where a language modeling head on top of the hidden states predicts the next tokens (?). Finally, the representational and generative objectives are summed with optional loss weights. Furthermore, sliding window attention (?, ?) is used by GRIT to handle generative and embedding inputs of arbitrary length.
The primary drawback of GRIT is its increased computational demand (as shown in Table 2), resulting from the need to train with two objective functions. GRITLM 7B is fine-tuned from Mistral 7B (?) and GRITLM 8x7B (?) is fine-tuned from Mistral 8x7B. Both models have top performance on MTEB English benchmark. GRITLM 7B has better performance than GRITLM 8X7B on embedding tasks, while GRITLM 8X7B is significantly better than GRITLM 7B on generative tasks. The authors provide some hypothesis on the reason why GRIT works on both embedding and generative tasks: 1. Generative language modeling and text embeddings are interconnected, requiring deep understanding of natural language, but expressed differently. 2. The unified model may contain parameters acting as a switch for either embedding tasks or generative tasks (?). Some other details about GRIT include:
  • Fine-tuning data: adaptations of E5 (?): adding S2ORC (?) to increase its scientific data (“E5S”); adaptations of Tülu 2 data (?): filtering out their custom prompts that contain answers related to the origin of their model.
@@
LLM based text embeddings have several disadvantages including high computational cost, longer response time and high embedding dimensions (which makes the downstream tasks training also computationally expensive). A recent paper named Versatile Text Embeddings Distilled from Large Language Models (?) tries to mitigate these limitations using knowledge distillation from LLMs with synthetic data generation and refinement, where queries are generated from LLMs given the contexts, and their positive and negative passages are mined and refined by LLMs.
The main contribution of (?) is designing the two-stage approach that uses LLMs to generate Few-shot Prompted Retrieval dataset (FRet). The first stage is LLM-based Diverse Query Generation: unlike (?), FRet uses LLM to analyze a selected web passage and produce both a description of the task $t$ and a pertinent query $q$ related to the task:
@@
where $p_{seed}$ is a passage drawn randomly from the web corpus $\mathcal{C}$ and $P_{QG}$ is a fixed few-shot prompt that is identical for every example. By drawing from a variety of free-form task descriptions, LLM is guided to generate a diverse set of queries. These pairs are subsequently utilized to train the embedding models, instructing the models to link a query and its associated instructions with the target passage (?). To further encourage the diversity of generated task descriptions and queries, many diverse task descriptions are added in the prompt.
@@
The second stage of FRet is LLM-based positive and negative mining. Unlike previous works (?, ?) assuming that the query $q$ generated from a given passage $p_{seed}$ forms a good training pair $(q,p_{seed})$, the authors hypothesize that there could be a better positive target passage for $q$ than $p_{seed}$ somewhere in the corpus of web passages as $p_{seed}$ is not guaranteed to maximize $P(p|q,t)$ over all the passages in the corpus (?). To mine better positives for the generated query, they train an initial embedding model using passage and generated query pairs $(q,p_{seed})$ with in-batch negatives. The trained embedding model is used to retrieve top $K$ neighbors $P=\{p^{(1)},...,p^{(K)}\}$ from the corpus given a generated query $q$. These retrieved passages are ranked by the LLM with two few-shot prompted LLM ranking functions:
@@
  • Query Likelihood (QL) (?): $QL(q,p)=LLM(q|p,\mathcal{P}_{QL})$, where $\mathcal{P}_{QL}$ is a prompt containing an instruction for judging query likelihood and several few-shot examples of relevant query and passage pairs (?).
@@
  • Relevance Classification (RC) (?): $RC(q,p)=LLM(label|p,\mathcal{P}_{RC})$, where $\mathcal{P}_{RC}$ is a prompt with few-shot examples for grading the relevance of each query-passage pair.
@@
The final ranking function $R(q,p)$ is obtained by combining the rankings from QL and RC with the standard Reciprocal Rank Fusion (RRF) approach (?). The top ranked passage is then selected as the new positive passage $p^{+}$ given the generated query ($p^{+}\neq p_{seed}$ happens for around 15% cases in their dataset). In terms of negative passage selection, they propose two methods: 1. a random nearest neighbor passage that is different from the original passage; 2. the k-th passage in the ranking. The generated FRet dataset has in total 6.6M examples, each containing a task, a query, a positive passage, and a negative passage (?). The authors propose a new embedding model Gecko based on a 1.2B parameter pre-trained transformer language model and fine-tuned on the generated FRet dataset, which is one of the top performing text embeddings on the English MTEB benchmark with small embedding dimensions (256 and 768). Some other details about Gecko include:
@@
  • Pre-training data: large-scale community QA dataset (?) with title-body text pairs from the Web.
@@
  • Unified fine-tuning data: FRet (6.6M) along with the following academic training datasets: Natural Questions (?), HotpotQA (?), FEVER (?), MedMCQA (?), SNLI (?), MNLI (?), and several classification datasets from Huggingface. For the multilingual model, training sets from MIRACL (?) is added.
@@
    • Fine-tuning: in-batch negative samples, hard negatives and same-tower negatives (other queries in the batch) (?)
  • Model size (google-gecko-preview-0409): 1.2B (backbone: gtr-t5-xl (?))
@@
The authors from GTE (?) also proposed their LLMs focused universal text embedding model based on Qwen1.5-7B large language model (?), which is one of the the top-ranking embedding models on both MTEB English and Chinese benchmarks. While the full details of fine-tuning are not disclosed, the authors summarized their key contributions as  [^8] https://huggingface.co/Alibaba-NLP/gte-Qwen1.5-7B-instruct: 1. the use of bidirectional attention mechanisms to enhance contextual understanding; 2. the use of instruction tuning; 3. the use of a large, multilingual text corpus that covers various domains and scenarios.
@@
In this section, the universal text embeddings leveraging LLMs (which make up the majority of the top 10 best performing models on the MTEB English benchmark) are introduced. Most of these models share the finding that LLMs acquire good text representations through comprehensive auto-regressive pre-training, requiring only minimal fine-tuning to become effective universal text embedding models. E5-mistral-7b-instruct from Microsoft and Gecko from Google DeepMind demonstrate two different ways to generate synthetic data from LLMs in order to improve universal text embeddings, while Echo-mistral (?) and LLM2Vec (?) show that good universal text embeddings can be achieved with the focus on enabling the bidirectional attentions for decoder only LLMs without using synthetic data. LoRA is widely used for the fine-tuning of LLMs based universal text embeddings, where LoRA ranks are found not affecting the overall performance substantially in (?). Instructions are used by all LLM focused text embedding models introduced in this section. One of the main reasons is the good instruction following ability of LLMs which makes them ideal for creating universal text embedding models that can handle a wide range of tasks using instructions. From Table 2, it can be told that Mistral-7B model is the most popular backbone model for LLMs focused text embeddings. One of the reasons is that enabling bidirectional attention (even without any training) works well for Mistral-7B. The authors from (?) speculate that Mistral models may be pre-trained with some form bidirectional attention. On the other hand, the full evaluation on MTEB of LLM based universal text embedding models is reported to be computationally expensive: it takes about 3 days on 8 V100 GPUs for E5-mistral-7b-instruct and 40 hours on 8x A100 GPUs for LLM2Vec with Mistral-7B as backbone.

### v2

@@ 5 LLMs focused universal text embeddings
LLMs are are extensively pre-trained on diverse large quantity of web-scale data, which can be used to improve the universal text embeddings in two different ways as summarized in Table 2. Firstly, LLMs can be used to generate high quality multilingual multi-task synthetic data as demonstrated by researchers from Microsoft and Google [16, 38, 34]. Secondly, LLMs can be used as backbone for text embeddings as they do not need the contrastive pre-training step used in existing state of the art text embedding models. For example, E5-mistral-7b-instruct perform multi-task fine-tuning on Mistral 7b model which is one of the best performing method on MTEB. Echo-mistral [48], LLM2Vec [39], gte-Qwen1.5-7B-instruct [12] and GRIT [49] propose various different solutions so that decoder only LLMs with casual attention can generate high quality text embeddings using bidirectional attention. More details about how different universal text embeddings leverage LLMs to improve their universality can be found below.
@@
E5-mistral-7b-instruct is one of the best performing text embeddings on the MTEB benchmark, which is also a representative text embedding model leveraging LLMs. Firstly, proprietary LLMs including GPT-35-Turbo and GPT-4 are used to generate synthetic data covering a diverse range of text embedding tasks in 93 languages (among which 25% are generated by GPT-35-Turbo and others are generated by GPT-4) [16]. In terms of the quality generated data, the authors find that the overall quality is acceptable despite a portion of GPT-35-Turbo outputs do not follow the instructions in the prompt templates strictly. Secondly, pre-trained open source LLM Mistral-7b checkpoint [61] is selected to be fine-tuned on a mixture of synthetic and labeled data (collection of 13 public datasets) with around 1.8M examples after sampling. One advantage of using LLMs such as Mistral [61] for text embedding is that they are extensively pre-trained on web-scale data already, which does not need the contrastive pre-training step used in existing state of the art text embedding models. Given a pre-trained LLM, an [EOS] token is appended to the end of the query and document. The last layer [EOS] vector is used as the text embeddings. To help the model accommodate different tasks, instruction templates (which are used by all LLMs focused universal text embeddings described in this section as well as some of the previously mentioned universal text embeddings such as BGE [35] ) are applied to the original query $q^{+}$ to generate a new one $q_{inst}^{+}$ given a relevant query-document pair $(q^{+},d^{+})$:
@@
where “task_definition” is a placeholder for a one-sentence description of the embedding task added only to the query side but not to the document side [16]. Some other details about E5-mistral-7b-instruct include:
  • Fine-tuning data: generated synthetic data, ELI5 [90] (sample ratio 0.1), HotpotQA [72], FEVER [75], MIRACL [91], MS-MARCO passage ranking (sample ratio 0.5) and document ranking (sample ratio 0.2) [69], NQ [70], NLI [74], SQuAD [70], TriviaQA [70], Quora Duplicate Questions [76] (sample ratio 0.1), Mr-TyDi [92], DuReader [93], and T2Ranking [94] (sample ratio 0.5) datasets.
@@
  • Negative sampling: in-batch negative samples and hard negatives (for the datasets without hard negatives, mE5base [38] is used to to mine top 100 hard negatives).
  • Model size: 7B (42M trainable parameters using Low-rank adaptation (LoRA) [95])
The experimental results from [16] shows that even with only synthetic data, the performance of E5-mistral-7b-instruct on MTEB English benchmark is still very competitive. E5-mistral-7b-instruct also has the multilingual capabilities with good performances over high-resource languages. Furthermore, the authors discovered that the method of incorporating instructions has a considerable impact on the performance. Their hypothesis is that the model is better informed about the embedding task at hand through natural language instructions, thereby allowing the model to produce more distinctive embeddings [16].
@@
Built on top of the E5-mistral-7b-instruct, SFR-Embedding-Mistral is also one of the top-ranking universal text embeddings on the MTEB English benchmark with 0.93% performance increase compared to E5-mistral-7b-instruct. The authors summarized their main takeaways in [51] (the detailed report is not released) as:
@@
Even though constructing text embeddings from autoregressive pretrained LLMs seems promising, the authors from [48] identified a striking failure mode of autoregressive language models trained on the next-token objective: the contextualized token embeddings, represented by the vector of last-hidden-layer activations at a specific input token’s position, lack information from tokens appearing later in the sentence because of the causal attention mask. Given the following example provided in [48]:
@@
In this example, the classical LLMs based contextualized embeddings of the first half of $d^{-}$ and $d^{+}$ are both similar to q because they do not attend to the second half of the sentence, which leads to the overestimation of the similarity between q and $d^{-}$ by any pooling strategy that uses information from the first half [48].
To mitigate this striking failure mode and take advantage of the bidirectional context information, a simple fix is proposed by presenting the input sentence twice to LLMs. The final contextualized embeddings can then be extracted from the second occurrence of the sentence. LLMs are instructed to undertake basic task such as rewriting or repeating in order to prompt the second occurrence to effectively ”encode” information from the first [48]. Despite twice the computational cost of classical embeddings, experimental results show that Echo embeddings can improve the LLM based text embedding quality significantly under both zero-shot setting and fine-tuning setting. Some other details about Echo-mistral (echo-mistral-7b-instruct-last) include:
  • Fine-tuning data: same as E5-mistral-7b-instruct [16] without synthetic data
@@
Similar to the idea of Echo-mistral, the authors of [39] believe that the slow adoption of decoder-only LLMs in text embedding tasks is partly due to their causal attention mechanism, which restricts their ability to create bidirectional contextualized representations from encompassing information from the whole input sequence (a necessary trade-off for generative capabilities). Improving the architectural flaw of decoder-only LLMs for text embedding tasks is highly desirable because: 1. decoder-only LLMs are much more sample-efficient than encoder-only models [96]; 2. LLMs are supported by a robust ecosystem, including comprehensive tools and well tested pre-training techniques, leading to their continuous enhancement by the community; 3. the good instruction following ability of LLMs [97, 98] makes them ideal for creating universal text embedding models that can handle a wide range of tasks using instructions.
To improve the text embeddings from decoder-only LLMs, LLM2Vec proposes a simple unsupervised approach that can transform any decoder-only LLM into a strong text encoder in three simple steps: 1. enabling bidirectional attention by replacing the causal attention mask of decoder-only LLMs with an all-ones matrix; 2. Masked Next Token Prediction (MNTP): combining next token prediction with masked language modeling [28] to make the model aware of its bidirectional attention; and 3. unsupervised contrastive learning for better sequence representations: the model processes an input sequence twice with independently sampled dropout masks to generate two distinct representations, and is trained to increase the similarity between these representations while decreasing similarity with other sequence representations in the batch [39]. Their empirical results show that LLMs can be efficiently converted into universal text embeddings without requiring costly adaptation or synthetic GPT-4 generated data. Some other details about LLM2Vec include:
@@
  • Supervised contrastive learning data: adaptations of E5 [16]: the public portion of the E5 dataset [16] curated by [38]
@@
Similar to the idea from Echo-mistral and LLM2Vec, the authors in [49] also highlight the importance of bidirectional attention for general purpose universal text embeddings. However, GRIT takes the general purpose model to the next level by training a large language model to handle both generative and embedding tasks (all text-based language problems) distinguished through instructions.
Both representational instruction tuning [77, 16, 36] and generative instruction tuning [99, 100, 101] are combined into one unified model by GRIT. Firstly, GRIT uses bidirectional attention with mean pooling over the final hidden state to get the text embedding. Contrastive objective with in-batch negatives are used to finetune a pretrained large language model following prior works [102, 57]. The average of the final hidden states of only the input sample is calculated during mean pooling, while disregarding the instruction and format tokens. Nonetheless, these tokens continue to impact the final representation via the self-attention mechanism [29]. Secondly, the language modeling objective of next token prediction [27] is used to compute the loss on generative data, where a language modeling head on top of the hidden states predicts the next tokens [49]. Finally, the representational and generative objectives are summed with optional loss weights. Furthermore, sliding window attention [103, 104] is used by GRIT to handle generative and embedding inputs of arbitrary length.
The primary drawback of GRIT is its increased computational demand (as shown in Table 2), resulting from the need to train with two objective functions. GRITLM 7B is fine-tuned from Mistral 7B [61] and GRITLM 8x7B [105] is fine-tuned from Mistral 8x7B. Both models have top performance on MTEB English benchmark. GRITLM 7B has better performance than GRITLM 8X7B on embedding tasks, while GRITLM 8X7B is significantly better than GRITLM 7B on generative tasks. The authors provide some hypothesis on the reason why GRIT works on both embedding and generative tasks: 1. Generative language modeling and text embeddings are interconnected, requiring deep understanding of natural language, but expressed differently. 2. The unified model may contain parameters acting as a switch for either embedding tasks or generative tasks [49]. Some other details about GRIT include:
  • Fine-tuning data: adaptations of E5 [16]: adding S2ORC [78] to increase its scientific data (“E5S”); adaptations of Tülu 2 data [106]: filtering out their custom prompts that contain answers related to the origin of their model.
@@
LLM based text embeddings have several disadvantages including high computational cost, longer response time and high embedding dimensions (which makes the downstream tasks training also computationally expensive). A recent paper named Versatile Text Embeddings Distilled from Large Language Models [34] tries to mitigate these limitations using knowledge distillation from LLMs with synthetic data generation and refinement, where queries are generated from LLMs given the contexts, and their positive and negative passages are mined and refined by LLMs.
The main contribution of [34] is designing the two-stage approach that uses LLMs to generate Few-shot Prompted Retrieval dataset (FRet). The first stage is LLM-based Diverse Query Generation: unlike [16], FRet uses LLM to analyze a selected web passage and produce both a description of the task $t$ and a pertinent query $q$ related to the task:
@@
where $p_{seed}$ is a passage drawn randomly from the web corpus $\mathcal{C}$ and $P_{QG}$ is a fixed few-shot prompt that is identical for every example. By drawing from a variety of free-form task descriptions, LLM is guided to generate a diverse set of queries. These pairs are subsequently utilized to train the embedding models, instructing the models to link a query and its associated instructions with the target passage [34]. To further encourage the diversity of generated task descriptions and queries, many diverse task descriptions are added in the prompt.
@@
The second stage of FRet is LLM-based positive and negative mining. Unlike previous works [107, 64] assuming that the query $q$ generated from a given passage $p_{seed}$ forms a good training pair $(q,p_{seed})$, the authors hypothesize that there could be a better positive target passage for $q$ than $p_{seed}$ somewhere in the corpus of web passages as $p_{seed}$ is not guaranteed to maximize $P(p|q,t)$ over all the passages in the corpus [39]. To mine better positives for the generated query, they train an initial embedding model using passage and generated query pairs $(q,p_{seed})$ with in-batch negatives. The trained embedding model is used to retrieve top $K$ neighbors $P=\{p^{(1)},...,p^{(K)}\}$ from the corpus given a generated query $q$. These retrieved passages are ranked by the LLM with two few-shot prompted LLM ranking functions:
@@
  • Query Likelihood (QL) [108]: $QL(q,p)=LLM(q|p,\mathcal{P}_{QL})$, where $\mathcal{P}_{QL}$ is a prompt containing an instruction for judging query likelihood and several few-shot examples of relevant query and passage pairs [109].
@@
  • Relevance Classification (RC) [110]: $RC(q,p)=LLM(label|p,\mathcal{P}_{RC})$, where $\mathcal{P}_{RC}$ is a prompt with few-shot examples for grading the relevance of each query-passage pair.
@@
The final ranking function $R(q,p)$ is obtained by combining the rankings from QL and RC with the standard Reciprocal Rank Fusion (RRF) approach [111]. The top ranked passage is then selected as the new positive passage $p^{+}$ given the generated query ($p^{+}\neq p_{seed}$ happens for around 15% cases in their dataset). In terms of negative passage selection, they propose two methods: 1. a random nearest neighbor passage that is different from the original passage; 2. the k-th passage in the ranking. The generated FRet dataset has in total 6.6M examples, each containing a task, a query, a positive passage, and a negative passage [34]. The authors propose a new embedding model Gecko based on a 1.2B parameter pre-trained transformer language model and fine-tuned on the generated FRet dataset, which is one of the top performing text embeddings on the English MTEB benchmark with small embedding dimensions (256 and 768). Some other details about Gecko include:
@@
  • Pre-training data: large-scale community QA dataset [112] with title-body text pairs from the Web.
@@
  • Unified fine-tuning data: FRet (6.6M) along with the following academic training datasets: Natural Questions [71], HotpotQA [72], FEVER [75], MedMCQA [113], SNLI [74], MNLI [73], and several classification datasets from Huggingface. For the multilingual model, training sets from MIRACL [91] is added.
@@
    • Fine-tuning: in-batch negative samples, hard negatives and same-tower negatives (other queries in the batch) [68]
  • Model size (google-gecko-preview-0409): 1.2B (backbone: gtr-t5-xl [112])
@@
The authors from GTE [12] also proposed their LLMs focused universal text embedding model based on Qwen1.5-7B large language model [114], which is one of the the top-ranking embedding models on both MTEB English and Chinese benchmarks. While the full details of fine-tuning are not disclosed, the authors summarized their key contributions as  [^8] https://huggingface.co/Alibaba-NLP/gte-Qwen1.5-7B-instruct: 1. the use of bidirectional attention mechanisms to enhance contextual understanding; 2. the use of instruction tuning; 3. the use of a large, multilingual text corpus that covers various domains and scenarios.
@@
In this section, the universal text embeddings leveraging LLMs (which make up the majority of the top 10 best performing models on the MTEB English benchmark) are introduced. Most of these models share the finding that LLMs acquire good text representations through comprehensive auto-regressive pre-training, requiring only minimal fine-tuning to become effective universal text embedding models. E5-mistral-7b-instruct from Microsoft and Gecko from Google DeepMind demonstrate two different ways to generate synthetic data from LLMs in order to improve universal text embeddings, while Echo-mistral [48] and LLM2Vec [39] show that good universal text embeddings can be achieved with the focus on enabling the bidirectional attentions for decoder only LLMs without using synthetic data. LoRA is widely used for the fine-tuning of LLMs based universal text embeddings, where LoRA ranks are found not affecting the overall performance substantially in [16]. Instructions are used by all LLM focused text embedding models introduced in this section. One of the main reasons is the good instruction following ability of LLMs which makes them ideal for creating universal text embedding models that can handle a wide range of tasks using instructions. From Table 2, it can be told that Mistral-7B model is the most popular backbone model for LLMs focused text embeddings. One of the reasons is that enabling bidirectional attention (even without any training) works well for Mistral-7B. The authors from [39] speculate that Mistral models may be pre-trained with some form bidirectional attention. On the other hand, the full evaluation on MTEB of LLM based universal text embedding models is reported to be computationally expensive: it takes about 3 days on 8 V100 GPUs for E5-mistral-7b-instruct and 40 hours on 8x A100 GPUs for LLM2Vec with Mistral-7B as backbone.